OpenStudy (anonymous):

How many amino acids could be specified if codons consisted of two nucleotides instead of three? What problem would this present for an organism making proteins using all 20 amino acids? MEDAL WILL BE GIVEN

OpenStudy (anonymous):

Codon combinations: A to T, C to G There are 4 possibilities: A to T or T to A making up 2 possibilities. C to G or G to C makes up the remaining possibilities. The necessary codons for 20 amino acids is simply the total number of amino acids 'times' the number of required condons: 20 amino acids times 3 codons = 60 codons. Thus, 4^2 ( 4 x 4 = 16), suggesting that codons based upon only 2 nucleotides would not be enough to code for the full range of 20 amino acids: 16 < 60. However, 4^3 ( 4 x 4 x 4 = 64) is enough to account for 20 amino acids; thus, 60 < 64 ------------------------------- Take the cube-root of 64: 64^(1/3) = 4 which is the number of possible A to T and C to G codon combinations.

OpenStudy (anonymous):

second question plz

OpenStudy (anonymous):

Go away =)

OpenStudy (anonymous):

And stop spamming